Solve by completing the square. 4x^2+2x-3=0
Answer by tanimachatterjee(60)

4X^2 + 2x - 3 = 0
or 4x^2 + 2. 2 x .1/2 + (1/2)^2 -(1/2)^2 -3 =0
or (2x + 1/2)^2 - (1/4 +3) =0
or ( 2x +1/2)^2 = 13/4
or 2x+1/2 = +- sqrt(13/4)
or 2x = +-sqrt 13/2 -1/2
or x = 1/2(+-sqrt 13/2 -1/2)
= 1/4(+- sqrt 13 -1)